Byron Books Inc. recently reported $6 million of net income. Its EBIT was $12.2 million, and its tax rate was 25%. What was its interest expense? (Hint: Write out the headings for an income statement, and then fill in the known values. Then divide $6 million of net income by (1 - T) = 0.75 to find the pretax income. The difference between EBIT and taxable income must be interest expense. Use this same procedure to complete similar problems.) Write out your answer completely. For example, 25 million should be entered as 25,000,000. Round your answer to the nearest dollar, if necessary. Do not round intermediate calculations.

$  
Answer #1

(EBIT-interest expense)(1-tax rate)=Net income

(12.2-interest expense)(1-0.25)=6

(12.2-interest expense)=6/0.75

interest expense=12.2-(6/0.75)

=$4.2 million

=$4,200,000